OpenStudy (anonymous):

this more of an accounting question but its mathy like if given the direct cost of sales=37000 and gross profit=75000 and Net profit=10%, what is the amount for the overhead expenses?

OpenStudy (anonymous):

somehow we are suppose to get 25,800

OpenStudy (amistre64):

cost = 37000 profit = 75000 gross is before taxes and net is after right?

OpenStudy (anonymous):

i believe so

OpenStudy (amistre64):

or is the 75000 spose to be revenue? and you have to minus the costs to get to the profit?

OpenStudy (amistre64):

saying that net profit is 10% and asking what overhead costs (fixed expenses perhaps) are seems rather odd

OpenStudy (anonymous):

i'm so confused about this stuff

OpenStudy (anonymous):

i don't what 10% means it doesn't say 10% of what

OpenStudy (amistre64):

net profit comes from gross profits; and taxes is the only thing that should affect it

OpenStudy (amistre64):

but if gross is 75000, then net would be 7500 to take home

OpenStudy (amistre64):

but thats beside the point; profit = revenue(money in) - costs(money out)

OpenStudy (anonymous):

37000-7500 is really close to 28500

OpenStudy (anonymous):

it only differs by 1000

OpenStudy (amistre64):

75000 = Revenue - 37000 - fixed costs

OpenStudy (anonymous):

wait no im wrong

OpenStudy (anonymous):

we want 25800 not 28500 :(

OpenStudy (amistre64):

this is more reasonably set up as: P(x) = 75000 - 37000 - fixed costs(F) P(x) = 38000 - F , but even then tere is not enough info that I see to make sense of the question

OpenStudy (amistre64):

re read the question and see if you have all the parts right or at least are not missing any part to it

OpenStudy (anonymous):

37000-75000(.1)-.1(37000)=25800

OpenStudy (anonymous):

37000-(.1)(75000+37000)

OpenStudy (anonymous):

Gross profit-netprofit%(Direct cost+gross profit) does this make sense?

OpenStudy (amistre64):

not according to any accounting classes ive taken ...

OpenStudy (anonymous):

lol we get 25800 from it

OpenStudy (amistre64):

if thats the answer you want, then go for it ;)

OpenStudy (anonymous):

i want one that make senses though

OpenStudy (anonymous):

just because it says it is the answer doesn't mean it is